OpenStudy (anonymous):

11d-9<(with line under)15d+3

OpenStudy (kirbykirby):

\[ 11d-9 \le 15d+3\\ 11d - 15d \le 3 + 9\\ -4d \le 12\\ \, \\ d \ge \frac{12}{-4}\\\, \\ d \ge -3\]

OpenStudy (kirbykirby):

switch the inequality's direction when you divide both sides by a negative number